THE ROLE OF πσ* STATE IN INTRAMOLECULAR CHARGE TRANSFER OF 4-(DIMETHYLAMINO)- BENZONITRILE AND RELATED MOLECULES

\underline{TAKASHIGE FUJIWARA}, Department of Physics, The Ohio State University, Columbus OH 43210; MAREK Z. ZGIERSKI, Steacie Institute for Molecular Science, National Research Council of Canada, Ottawa, K1A 0R6 CANADA; EDWARD C. LIM, Department of Chemistry and The Center for Laser and Optical Spectroscopy, The University of Akron, Akron OH 44325-3601.

The solvent-polarity dependence and temporal characteristics of the transient absorption of 4-(dimethylamino)benzonitrile, DMABN, and 4-(dimethylamino)benzethyne, DMABE, demonstrate the presence of the πσ*-state absorption at about 700 nm and the ππ* (LE)- state absorption at about 520 nm and 450 nm. The rise and decay times of the πσ*-state transient differ from those of the ππ*-state transients in both compounds. Moreover, the peak position of the πσ*-state absorption is blue-shifted and more intense in acetonitrile as compared to n-hexane, whereas the band positions of the ππ*-state absorptions are essentially the same in the two solvents. For DMABN in acetonitrile, the rise time (∼ 4.3 ps) of the twisted intramolecular charge transfer (TICT)-state transient at 330 nm is identical to the decay time of the πσ*-state transient. The 4.8 ns decay time of the TICT-state absorption of DMABN is longer than the 2.9 ns decay time of the intramolecular charge-transfer (ICT) fluorescence, indicating that the fluorescent ICT state differs from the TICT state observed in transient absorption. These results are consistent with the presence of a low-lying πσ* state in DMABN (and DMABE), and the role the πσ* state plays in the formation of the TICT state of DMABN.